# Don’t Forget Its Variance! The Minimum Path Variance Principle for Accurate and Stable Score-Based Models

### v2

@@ Abstract
Score-based methods are powerful across machine learning, but they face a paradox: theoretically path-independent, yet practically path-dependent. We resolve this by proving that practical training objectives differ from the ideal, ground-truth objective by a crucial, overlooked term: the path variance of the score function. We propose the MinPV (Minimum Path Variance) Principle to minimize this path variance. Our key contribution is deriving a closed-form expression for the variance, making optimization tractable. By parameterizing the path with a flexible Kumaraswamy Mixture Model, our method learns data-adaptive, low-variance paths without heuristic manual selection. This principled optimization of the complete objective yields more accurate and stable estimators, establishing new state-of-the-art results on challenging benchmarks and providing a general framework for optimizing score-based interpolation.
@@ 1 Introduction
Our main contribution is to close this gap. We prove that the missing term is precisely the path variance of the ground-truth score, a quantity that must be minimized and cannot be treated as a constant (illustrated in Figs. 1(b) and 1(c) and proved in Sec. 4.1). Building on this insight, we propose MinPV (Minimum Path Variance) Principle, the first framework that directly optimize the complete, ideal objective. Our contributions are threefold:
@@ 2 Related Works
We propose the first DRE framework that goes beyond fixed or heuristic paths by explicitly learning an optimal path through a theoretically grounded objective. Our analysis shows that this objective (the path variance) is the missing term in the ideal score-matching loss that prior DRE methods have overlooked. By parameterizing the path with a flexible KMM and optimizing it with its path variance, we convert the intractable functional search into a tractable, low-dimensional optimization. Our method, MinPV, thus adapts the path schedule to the data distributions, achieving more accurate density ratio estimation.
@@ 3 Background
To approximate $s^{(t)}$, we train a time score model $s_{\bm{\theta}}^{(t)}$ by minimizing the ideal but intractable time score matching objective, $\mathcal{L}_{\text{TSM}}({\bm{\theta}})=\mathbb{E}_{p(t)p_{t}({\bm{x}})}\left|s^{(t)}({\bm{x}},t)-s_{{\bm{\theta}}}^{(t)}({\bm{x}},t)\right|^{2}$. Since $s^{(t)}$ is unknown, this loss is replaced by a tractable alternative derived via integration by parts, known as the sliced time score matching (STSM)  [^1] We adopt the term “sliced” because the objective, derived through integration by parts, involves projecting onto random directions, a technique known as Sliced Score Matching (Song et al., 2020b) in the generative modeling literature. Choi et al. (2022) simply call this objective Time Score Matching. (Choi et al., 2022) objective,
@@ 4 Minimum Path Variance Principle
Our proposed method, MinPV, is designed to resolve an important discrepancy between the theory and practice of score-based DRE. In this section, we detail the core components of our method: the derivation of the analytical path variance and our flexible KMM-based path parameterization.
@@ 4.1 The Path Variance: An Overlooked Term in the Objective
This establishes the Minimum Path Variance (MinPV) principle: Minimizing the overall error bound requires jointly minimizing the tractable model loss $\mathcal{L}_{\textnormal{STSM}}$ and the path variance $\mathcal{V}$. See Appendix Sec. A.3 for proof. This theorem reveals that path variance is not an auxiliary regularizer but the overlooked component of the ideal objective. Conventional approaches fix the path a priori, treating $\mathcal{V}$ as constant and ignoring its effect. Our key insight is that, by learning the path itself, we can explicitly minimize this term, not as a guarantee of a tighter bound, but as a heuristic that empirically improves estimation accuracy and stability.
@@ 4.2 Minimizing the Path Variance via KMM Parameterization
Algorithm 1 DRE with MinPV: Inference.
@@
The path parameters $\bm{\phi}$} and score model parameters $\bm{\theta}$} are jointly optimized via the total objective $\mathcal{L}_{\text{total}}=\lambda_{1}\mathcal{L}_{\text{STSM}}({\bm{\theta}})+\lambda_{2}\mathcal{V}[\alpha_{{\bm{\phi}}},\beta_{{\bm{\phi}}}]$ (described in Appendix Sec. B.2). Details of path optimization are given in Algorithm 2, and a PyTorch implementation is provided in Algorithm 3 (see appendix). After training, MinPV uses the same inference as conventional DRE methods shown in Algorithm 1.
Algorithm 2 Unconstrained Optimization of the MinPV Path.
@@ 5 Experiments
We conduct a comprehensive set of experiments to validate the effectiveness of our Minimum Path Variance principle and demonstrate the superior performance of our proposed method, MinPV.
@@ 5.1 Baseline Setting and Experimental Setup
To evaluate our approach, we benchmark against a comprehensive suite of common, fixed path schedules representing the current standard in the DRE and generative modeling literatures. These baselines include the Linear (Albergo et al., 2023), VP (Ho et al., 2020; Song et al., 2020a), Cosine (Nichol and Dhariwal, 2021), Föllmer (Föllmer, 2005; Chen et al., 2024), and Trigonometric (Albergo et al., 2023) paths, with detailed formulations provided in Sec. B.1. In contrast to these pre-defined paths, our MinPV path schedule utilizes a learnable path parameterized by a KMM, as discussed in Sec. 4.2 above. The path is optimized by directly minimizing the analytical path variance expressions from Sec. 4.1, allowing it to adapt to the specific data distributions of a given task.
@@
We first conduct an ablation study on the sensitivity of MinPV to the number of components $K$ in the KMM path parameterization. Experiments are done under the most challenging settings for both tasks: mutual information (MI) estimation at $d=160$ and density estimation on BSDS300. As shown in Tab. 1, a single component ($K=1$) is too restrictive and performs poorly, highlighting the need for path flexibility. Performance improves markedly at $K=2$ and reaches its peak around $K=5$. Increasing the number of components further ($K=8$) brings no benefit and may slightly hurt performance due to overfitting the path variance objective. Thus, MinPV empirically remains robust, with $K\in[2,5]$. In practice, we set $K=5$, which achieves the best trade-off.
@@ 5.3 $f$-divergence Estimation
We first test robustness on five mutual information (MI) estimation tasks that can be viewed as challenging $f$-divergence problems. (For a formal MI definition, see Appendix Sec. B.4.) Following Czyż et al. (2023), these distributions exhibit concentrated densities, sharp discontinuities, heavy tails, and extreme correlations, all of which break standard estimator assumptions. Results in Tab. 2 show that fixed-path baselines degrade severely on these geometries, while MinPV consistently achieves low error (see also Tab. 6(e) in the appendix). On the Additive Noise task with discontinuities and the Gamma–Exponential task with complex non-linear dependence, our method outperforms all baselines by a wide margin, confirming that path-variance optimization enables MinPV to adapt to pathological structures effectively.
Table 2: MSE results on the Additive Noise (sharp discontinuities) and Gamma–Exponential (non-linear dependency) datasets. Across all correlation levels (top row of each sub-table), MinPV achieves notably lower MSE than fixed-path baselines. Full results on 5 datasets given in Tab. 6(e) in appendix.
@@
MinPV (ours) affine $0.2360$ $0.0199$ $0.0081$ $0.0167$ $0.0161$ $0.0216$ $0.0304$ $0.0155$ $\mathbf{0.0010}$
MinPV (ours) spherical $\mathbf{0.1016}$ $\mathbf{0.0134}$ $\mathbf{0.0078}$ $\mathbf{0.0009}$ $\mathbf{0.0138}$ $\mathbf{0.0117}$ $\mathbf{0.0196}$ $\mathbf{0.0026}$ $0.0029$
@@
MinPV (ours) affine $0.4785$ $1.5466$ $0.1137$ $0.1274$ $0.0915$ $\mathbf{0.0063}$ $\mathbf{0.0032}$ $0.0294$ $0.0069$
MinPV (ours) spherical $\mathbf{0.2721}$ $\mathbf{0.2304}$ $\mathbf{0.0917}$ $\mathbf{0.0434}$ $\mathbf{0.0075}$ $0.0187$ $0.0091$ $\mathbf{0.0030}$ $\mathbf{0.0004}$
@@
We further evaluate MI estimation in a high-dimensional, high-discrepancy setting, designed to induce the “density-chasm” problem in DRE (Rhodes et al., 2020). Performance is measured by MI accuracy and MSE against the true value. As shown in Tab. 3, most fixed-path baselines collapse as dimension $d$ and MI increase. For example, Föllmer and Cosine paths perform reasonably at low $d$ but diverge at $d=160$, MI=40=40. In contrast, MinPV with an affine constraint maintains a low MSE of $1.02$, outperforming the next best baseline by a large margin. These results highlight that adaptive path learning under our Minmimum Path Variance principle reliably bridges the density chasm in settings where heuristic fixed paths fail.
Table 3: Mutual information estimation under high-discrepancy settings (MI $\in\{10,20,30,40\}$ nats). We report the estimated mutual information (mean $\pm$ std) and MSE across different path settings and constraint types. Bolded MSE values indicate the best performance for each dimension. Our MinPV path demonstrates superior performance in most high-discrepancy settings.
@@
MinPV (ours) Spherical $9.81\scriptstyle\pm 0.24$ $0.16$ $20.43\scriptstyle\pm 0.44$ $0.62$ $28.88\scriptstyle\pm 0.69$ $2.35$ $37.22\scriptstyle\pm 0.72$ $8.79$
MinPV (ours) Affine $8\scriptstyle\pm 0.20 0.10\mathbf{0.10}$ $19.80\scriptstyle\pm 0.36$ $\mathbf{0.34}$ $30.22\scriptstyle\pm 0.48$ $\mathbf{0.58}$ $40.59\scriptstyle\pm 0.55$ $\mathbf{1.02}$
@@ 5.4 Density Estimation
Figure 2: Density estimation on checkerboard and tree datasets. MinPV successfully learns a data-adaptive path tailored to the specific manifold of each dataset. Full results in Fig. 4 of appendix.
@@
We further evaluate MinPV on six challenging datasets: circles, rings, pinwheel, 2spirals, checkerboard, and tree (Grathwohl et al., 2018; Karras et al., 2024), which feature disconnected components, intricate spirals, and discontinuous densities. As shown in Fig. 2, MinPV yields sharper and more accurate density estimates on checkerboard and tree, and matches the best baseline (VP path) on the remaining datasets.
Full results are provided in Fig. 4 in the appendix. These results demonstrate the significant adaptability of MinPV and help confirm that minimizing path variance enables MinPV to adapt effectively to diverse data geometries without manual path selection.
@@
Finally, we evaluate MinPV on five widely-used tabular benchmarks (Grathwohl et al., 2018), reporting test Negative Log-Likelihood (NLL), where lower is better. As shown in Tab. 4, MinPV consistently outperforms all fixed-path baselines and achieves new state-of-the-art results across all datasets. Two key observations of note are: 1) On POWER and GAS, the affine constraint leads to the best performance, whereas the spherical constraint is more effective on the higher-dimensional datasets (HEPMASS, MINIBOONE, and BSDS300). This highlights the adaptability of MinPV to data geometry. 2) The gains over strong baselines (e.g., VP and trigonometric paths) are substantial, especially on BSDS300 where MinPV improves NLL by more than 10 points, underscoring the advantage of explicitly minimizing path variance. These results validate our central claim, that adaptively selecting the path via MinPV yields a data-dependent path that aligns better with the underlying structure, leading to improved density estimation.
Table 4: Test Negative Log-Likelihood (NLL) on five tabular datasets (lower is better). Our MinPV path consistently outperforms all fixed-path baselines, achieving SOTA across the benchmarks.
@@
MinPV (ours) Spherical $0.15\scriptstyle\scriptstyle\pm 0.11$ $6.92\scriptstyle\scriptstyle\pm 0.21$ $\mathbf{18.01}\scriptstyle\scriptstyle\pm 0.16$ $\mathbf{17.81}\scriptstyle\scriptstyle\pm 0.09$ $\mathbf{-143.97}\scriptstyle\scriptstyle\pm 0.22$
MinPV (ours) Affine $\mathbf{-0.81}\scriptstyle\scriptstyle\pm 0.08$ $\mathbf{-6.95}\scriptstyle\scriptstyle\pm 0.10$ $18.79\scriptstyle\scriptstyle\pm 0.16$ $17.99\scriptstyle\scriptstyle\pm 0.10$ $129.73\scriptstyle\scriptstyle\pm 0.20$
@@ 5.5 Illustration of Optimized Path Schedules
We visualize the optimized schedules $\alpha_{{\bm{\phi}}}(t)$ and $\beta_{{\bm{\phi}}}(t)$ across various tasks to interpret the efficacy of the MinPV principle, with detailed comparisons provided in Fig. 3 and Sec. B.6. While Sec. 4.1 establishes a theoretical link between path variance and estimation error , we emphasize that our implementation treats the minimization of $\mathcal{V}$ as a principled heuristic rather than a strict minimization of the ground-truth error. This analytical proxy objective serves to encourage smoother path transitions and indirectly control the Lipschitz constant $L$. As illustrated in the optimized schedules for pathological and high-dimensional distributions Figs. 5 and 6, the learned paths (red) significantly deviate from fixed baselines. Specifically, MinPV tends to adopt a more gradual rate of change near the boundaries $t\in\{0,1\}$, which serves as a heuristic to suppress instantaneous velocity spikes in the time score that often lead to numerical instability.
@@
This data-adaptive schedule is particularly effective for navigating the density-chasm encountered in high-dimensional settings. As shown in Figs. 3 and 6, while fixed schedules like Föllmer or cosine often traverse regions of vanishingly low density and subsequently diverge, the MinPV path adapts its velocity to maintain more reliable score signals. The structural diversity of the learned schedules across structured manifolds Fig. 7 and real-world tabular datasets Fig. 8 further underscores the necessity of the flexible KMM parameterization. By allowing the schedule to adjust its rate of change to align with the specific manifold structure, MinPV ensures that the probability flow is tailored to the data distribution. This alignment directly facilitates the substantial improvements in NLL and SOTA results reported in Tab. 4.

### v3

# A Minimum Variance Path Principle for Accurate and Stable Score-Based Density Ratio Estimation

@@ Abstract
Score-based methods are powerful across machine learning, but they face a paradox: theoretically path-independent, yet practically path-dependent. We resolve this by proving that practical training objectives differ from the ideal, ground-truth objective by a crucial, overlooked term: the path variance of the score function. We propose the MVP (Mimum Variance Path) Principle to minimize this path variance. Our key contribution is deriving a closed-form expression for the variance, making optimization tractable. By parameterizing the path with a flexible Kumaraswamy Mixture Model, our method learns data-adaptive, low-variance paths without heuristic manual selection. This principled optimization of the complete objective yields more accurate and stable estimators, establishing new state-of-the-art results on challenging benchmarks and providing a general framework for optimizing score-based interpolation.
@@ 1 Introduction
Our main contribution is to close this gap. We prove that the missing term is precisely the path variance of the ground-truth score, a quantity that must be minimized and cannot be treated as a constant (illustrated in Figs. 1(b) and 1(c) and proved in Sec. 4.1). Building on this insight, we propose MVP (Mimum Variance Path) Principle, the first framework that directly optimize the complete, ideal objective. Our contributions are threefold:
@@ 2 Related Works
We propose the first DRE framework that goes beyond fixed or heuristic paths by explicitly learning an optimal path through a theoretically grounded objective. Our analysis shows that this objective (the path variance) is the missing term in the ideal score-matching loss that prior DRE methods have overlooked. By parameterizing the path with a flexible KMM and optimizing it with its path variance, we convert the intractable functional search into a tractable, low-dimensional optimization. Our method, MVP, thus adapts the path schedule to the data distributions, achieving more accurate density ratio estimation.
@@ 3 Background
To approximate $s^{(t)}$, we train a time score model $s_{\bm{\theta}}^{(t)}$ by minimizing the ideal but intractable time score matching objective, $\mathcal{L}_{\text{TSM}}({\bm{\theta}})=\mathbb{E}_{p(t)p_{t}({\bm{x}})}\left|s^{(t)}({\bm{x}},t)-s_{{\bm{\theta}}}^{(t)}({\bm{x}},t)\right|^{2}$. Since $s^{(t)}$ is unknown, this loss is replaced by a tractable alternative derived via integration by parts, known as the sliced time score matching (STSM)  [^1] We adopt the term “sliced” because the objective, derived through integration by parts, involves projecting onto random directions, a technique known as Sliced Score Matching (Song et al., 2020) in the generative modeling literature. Choi et al. (2022) simply call this objective Time Score Matching. (Choi et al., 2022) objective,
@@ 4 Minimum Path Variance Principle
Our proposed method, MVP, is designed to resolve an important discrepancy between the theory and practice of score-based DRE. In this section, we detail the core components of our method: the derivation of the analytical path variance and our flexible KMM-based path parameterization.
@@ 4.1 The Path Variance: An Overlooked Term in the Objective
This establishes the Minimum Variance Path (MVP) principle: Minimizing the overall error bound requires jointly minimizing the tractable model loss $\mathcal{L}_{\textnormal{STSM}}$ and the path variance $\mathcal{V}$. See Appendix Sec. A.3 for proof. This theorem reveals that path variance is not an auxiliary regularizer but the overlooked component of the ideal objective. Conventional approaches fix the path a priori, treating $\mathcal{V}$ as constant and ignoring its effect. Our key insight is that, by learning the path itself, we can explicitly minimize this term, not as a guarantee of a tighter bound, but as a heuristic that empirically improves estimation accuracy and stability.
@@ 4.2 Minimizing the Path Variance via KMM Parameterization
Algorithm 1 DRE with MVP: Inference.
@@
The path parameters $\bm{\phi}$} and score model parameters $\bm{\theta}$} are jointly optimized via the total objective $\mathcal{L}_{\text{total}}=\lambda_{1}\mathcal{L}_{\text{STSM}}({\bm{\theta}})+\lambda_{2}\mathcal{V}[\alpha_{{\bm{\phi}}},\beta_{{\bm{\phi}}}]$ (described in Appendix Sec. B.2). Details of path optimization are given in Algorithm 2, and a PyTorch implementation is provided in Algorithm 3 (see appendix). After training, MVP uses the same inference as conventional DRE methods shown in Algorithm 1.
Algorithm 2 Unconstrained Optimization of the MVP Path.
@@ 5 Experiments
We conduct a comprehensive set of experiments to validate the effectiveness of our Minimum Path Variance principle and demonstrate the superior performance of our proposed method, MVP.
@@ 5.1 Baseline Setting and Experimental Setup
To evaluate our approach, we benchmark against a comprehensive suite of common, fixed path schedules representing the current standard in the DRE and generative modeling literatures. These baselines include the Linear (Albergo et al., 2023), VP (Ho et al., 2020; Song et al., 2021b), Cosine (Nichol and Dhariwal, 2021), Föllmer (Föllmer, 2005; Chen et al., 2024), and Trigonometric (Albergo et al., 2023) paths, with detailed formulations provided in Sec. B.1. In contrast to these pre-defined paths, our MVP path schedule utilizes a learnable path parameterized by a KMM, as discussed in Sec. 4.2 above. The path is optimized by directly minimizing the analytical path variance expressions from Sec. 4.1, allowing it to adapt to the specific data distributions of a given task.
@@
We first conduct an ablation study on the sensitivity of MVP to the number of components $K$ in the KMM path parameterization. Experiments are done under the most challenging settings for both tasks: mutual information (MI) estimation at $d=160$ and density estimation on BSDS300. As shown in Tab. 1, a single component ($K=1$) is too restrictive and performs poorly, highlighting the need for path flexibility. Performance improves markedly at $K=2$ and reaches its peak around $K=5$. Increasing the number of components further ($K=8$) brings no benefit and may slightly hurt performance due to overfitting the path variance objective. Thus, MVP empirically remains robust, with $K\in[2,5]$. In practice, we set $K=5$, which achieves the best trade-off.
@@ 5.3 $f$-divergence Estimation
We first test robustness on five mutual information (MI) estimation tasks that can be viewed as challenging $f$-divergence problems. (For a formal MI definition, see Appendix Sec. B.4.) Following Czyż et al. (2023), these distributions exhibit concentrated densities, sharp discontinuities, heavy tails, and extreme correlations, all of which break standard estimator assumptions. Results in Tab. 2 show that fixed-path baselines degrade severely on these geometries, while MVP consistently achieves low error (see also Tab. 6(e) in the appendix). On the Additive Noise task with discontinuities and the Gamma–Exponential task with complex non-linear dependence, our method outperforms all baselines by a wide margin, confirming that path-variance optimization enables MVP to adapt to pathological structures effectively.
Table 2: MSE results on the Additive Noise (sharp discontinuities) and Gamma–Exponential (non-linear dependency) datasets. Across all correlation levels (top row of each sub-table), MVP achieves notably lower MSE than fixed-path baselines. Full results on 5 datasets given in Tab. 6(e) in appendix.
@@
MVP (ours) affine $0.2360$ $0.0199$ $0.0081$ $0.0167$ $0.0161$ $0.0216$ $0.0304$ $0.0155$ $\mathbf{0.0010}$
MVP (ours) spherical $\mathbf{0.1016}$ $\mathbf{0.0134}$ $\mathbf{0.0078}$ $\mathbf{0.0009}$ $\mathbf{0.0138}$ $\mathbf{0.0117}$ $\mathbf{0.0196}$ $\mathbf{0.0026}$ $0.0029$
@@
MVP (ours) affine $0.4785$ $1.5466$ $0.1137$ $0.1274$ $0.0915$ $\mathbf{0.0063}$ $\mathbf{0.0032}$ $0.0294$ $0.0069$
MVP (ours) spherical $\mathbf{0.2721}$ $\mathbf{0.2304}$ $\mathbf{0.0917}$ $\mathbf{0.0434}$ $\mathbf{0.0075}$ $0.0187$ $0.0091$ $\mathbf{0.0030}$ $\mathbf{0.0004}$
@@
We further evaluate MI estimation in a high-dimensional, high-discrepancy setting, designed to induce the “density-chasm” problem in DRE (Rhodes et al., 2020). Performance is measured by MI accuracy and MSE against the true value. As shown in Tab. 3, most fixed-path baselines collapse as dimension $d$ and MI increase. For example, Föllmer and Cosine paths perform reasonably at low $d$ but diverge at $d=160$, MI=40=40. In contrast, MVP with an affine constraint maintains a low MSE of $1.02$, outperforming the next best baseline by a large margin. These results highlight that adaptive path learning under our Minmimum Path Variance principle reliably bridges the density chasm in settings where heuristic fixed paths fail.
Table 3: Mutual information estimation under high-discrepancy settings (MI $\in\{10,20,30,40\}$ nats). We report the estimated mutual information (mean $\pm$ std) and MSE across different path settings and constraint types. Bolded MSE values indicate the best performance for each dimension. Our MVP path demonstrates superior performance in most high-discrepancy settings.
@@
MVP (ours) Spherical $9.81\scriptstyle\pm 0.24$ $0.16$ $20.43\scriptstyle\pm 0.44$ $0.62$ $28.88\scriptstyle\pm 0.69$ $2.35$ $37.22\scriptstyle\pm 0.72$ $8.79$
MVP (ours) Affine $8\scriptstyle\pm 0.20 0.10\mathbf{0.10}$ $19.80\scriptstyle\pm 0.36$ $\mathbf{0.34}$ $30.22\scriptstyle\pm 0.48$ $\mathbf{0.58}$ $40.59\scriptstyle\pm 0.55$ $\mathbf{1.02}$
@@ 5.4 Density Estimation
Figure 2: Density estimation on checkerboard and tree datasets. MVP successfully learns a data-adaptive path tailored to the specific manifold of each dataset. Full results in Fig. 4 of appendix.
@@
We further evaluate MVP on six challenging datasets: circles, rings, pinwheel, 2spirals, checkerboard, and tree (Grathwohl et al., 2018; Karras et al., 2024), which feature disconnected components, intricate spirals, and discontinuous densities. As shown in Fig. 2, MVP yields sharper and more accurate density estimates on checkerboard and tree, and matches the best baseline (VP path) on the remaining datasets.
Full results are provided in Fig. 4 in the appendix. These results demonstrate the significant adaptability of MVP and help confirm that minimizing path variance enables MVP to adapt effectively to diverse data geometries without manual path selection.
@@
Finally, we evaluate MVP on five widely-used tabular benchmarks (Grathwohl et al., 2018), reporting test Negative Log-Likelihood (NLL), where lower is better. As shown in Tab. 4, MVP consistently outperforms all fixed-path baselines and achieves new state-of-the-art results across all datasets. Two key observations of note are: 1) On POWER and GAS, the affine constraint leads to the best performance, whereas the spherical constraint is more effective on the higher-dimensional datasets (HEPMASS, MINIBOONE, and BSDS300). This highlights the adaptability of MVP to data geometry. 2) The gains over strong baselines (e.g., VP and trigonometric paths) are substantial, especially on BSDS300 where MVP improves NLL by more than 10 points, underscoring the advantage of explicitly minimizing path variance. These results validate our central claim, that adaptively selecting the path via MVP yields a data-dependent path that aligns better with the underlying structure, leading to improved density estimation.
Table 4: Test Negative Log-Likelihood (NLL) on five tabular datasets (lower is better). Our MVP path consistently outperforms all fixed-path baselines, achieving SOTA across the benchmarks.
@@
MVP (ours) Spherical $0.15\scriptstyle\scriptstyle\pm 0.11$ $6.92\scriptstyle\scriptstyle\pm 0.21$ $\mathbf{18.01}\scriptstyle\scriptstyle\pm 0.16$ $\mathbf{17.81}\scriptstyle\scriptstyle\pm 0.09$ $\mathbf{-143.97}\scriptstyle\scriptstyle\pm 0.22$
MVP (ours) Affine $\mathbf{-0.81}\scriptstyle\scriptstyle\pm 0.08$ $\mathbf{-6.95}\scriptstyle\scriptstyle\pm 0.10$ $18.79\scriptstyle\scriptstyle\pm 0.16$ $17.99\scriptstyle\scriptstyle\pm 0.10$ $129.73\scriptstyle\scriptstyle\pm 0.20$
@@ 5.5 Illustration of Optimized Path Schedules
We visualize the optimized schedules $\alpha_{{\bm{\phi}}}(t)$ and $\beta_{{\bm{\phi}}}(t)$ across various tasks to interpret the efficacy of the MVP principle, with detailed comparisons provided in Fig. 3 and Sec. B.6. While Sec. 4.1 establishes a theoretical link between path variance and estimation error , we emphasize that our implementation treats the minimization of $\mathcal{V}$ as a principled heuristic rather than a strict minimization of the ground-truth error. This analytical proxy objective serves to encourage smoother path transitions and indirectly control the Lipschitz constant $L$. As illustrated in the optimized schedules for pathological and high-dimensional distributions Figs. 5 and 6, the learned paths (red) significantly deviate from fixed baselines. Specifically, MVP tends to adopt a more gradual rate of change near the boundaries $t\in\{0,1\}$, which serves as a heuristic to suppress instantaneous velocity spikes in the time score that often lead to numerical instability.
@@
This data-adaptive schedule is particularly effective for navigating the density-chasm encountered in high-dimensional settings. As shown in Figs. 3 and 6, while fixed schedules like Föllmer or cosine often traverse regions of vanishingly low density and subsequently diverge, the MVP path adapts its velocity to maintain more reliable score signals. The structural diversity of the learned schedules across structured manifolds Fig. 7 and real-world tabular datasets Fig. 8 further underscores the necessity of the flexible KMM parameterization. By allowing the schedule to adjust its rate of change to align with the specific manifold structure, MVP ensures that the probability flow is tailored to the data distribution. This alignment directly facilitates the substantial improvements in NLL and SOTA results reported in Tab. 4.